# Can Code-Switched Texts Activate a Knowledge Switch in LLMs? A Case Study on English-Korean Code-Switching

### v1

@@ 4.2 Experimental Settings
We conduct experiments using the same models as in Section 3.2. For knowledge identification, we instruct the model to write a list of factual knowledge that are required for solving the given question in one or two sentences. For knowledge leveraging, we pass on previously identified knowledge and ask the model to select an answer and explain why. Additionally, we inform that a correct answer exists in one of the choices, and request the not to respond “none of the above”. The full-length prompt for Knowledge Identification is provided in Table 6, and the prompt for Knowledge Leveraging is detailed in Table 7.
5 Results
In this section, we discuss the experimental results and analysis on the tasks from Section 4.
5.1 Analysis on Knowledge Identification
@@
In the upper row of Figure 4, we observe a significant gap in faithfulness scores between CS and English in both the History and Tradition domains. The discrepancy is more salient in Tradition where cultural nuances is much important, implying that asking questions in CS is much successful in capturing cultural nuances and meanings. In the General domain, the scores for CS and English are almost identical (or even better in English for Gemma2 9B), indicating that the difference in knowledge activated by CS questions compared to English questions is minimal when addressing general and common facts. In the Law domain, although knowledge from CS is slightly more faithful than that from English, their absolute scores are lower than those in other domains. As the Law domain requires expert knowledge, it is visible that models fail to identify faithful knowledge.
@@
The lower row of Figure 4 presents evaluation results for helpfulness. It is intuitive that faithful knowledge serves as a valuable source for answering questions, and as a result, the evaluation of helpfulness shows a similar trend to that of faithfulness. In the History and Tradition domains, the gap between CS and English becomes larger in helpfulness, emphasizing the effectiveness of the CS setting in identifying both faithful and helpful knowledge. It is also notable that the scores for helpfulness are particularly high for GPT-4o and Solar, models in which performance in CS surpasses that in the English setting to a large extent (§ 3.3). In contrast, the helpfulness scores in the Law domain are considerably lower for both CS and English compared to other domains. Given that the Law domain requires expert-level legal knowledge, the models struggle to grasp the legal context, leading to difficulties in accurately identifying helpful knowledge sources from both CS and English questions.
@@
In Figure 5, the win ratio for CS is higher in the History and Tradition domains, demonstrating that CS questions can activate more essential knowledge sources for question answering. On the contrary, in domains where CS does not show its effectiveness, the win ratio of CS is comparatively lower (i.e., General) or the ratio of Tie is high (i.e., Law). Especially in the Law domain, the quality of knowledge lists generated from CS questions is evaluated as equivalent to, or even worse than, that generated from English questions.
5.2 Analysis on Knowledge Leveraging
@@
We present the visualized results of QA accuracy in both CS and English settings in Figure 6, with detailed scores reported in Table 8. Consistent with the results in Section 3.3, all models demonstrate generally higher accuracy for CS questions compared to English questions, indicating that CS effectively activates knowledge across various domains. To be specific, the performance in the CS setting exceeds that of English in every domain for GPT-4o, Claude 3.5, and Solar. These models not only identify faithful and helpful knowledge (§ 5.1), but also answer questions while accurately grounding on that knowledge; this shows that CS questions robustly activate essential knowledge. Notably, these models show excellent performance on Korean (Figure 2), strengthening the correlation between monolingual (i.e., Korean) and CS knowledge activation abilities. On the contrary, Llama3 and Gemma2 families do not seem to benefit from CS questions. Specifically, these models show poor performance in both CS and English settings in several domains, such as Geography and Law. Taking into account that these domains require domain-specific expertise, it is likely that their lack of understanding contributes to low accuracy, let alone CS failing to activate Korea-focused knowledge.
5.3 Case Study
We examine sample cases to compare the capability of CS on knowledge activation over English. Table 10 shows the knowledge and answer generated by Solar in the Tradition domain. The question asks about “정월대보름”, a Korean traditional holiday that celebrates the first full moon of lunar new year. It is observable that CS question preserves unique terms such as “정월대보름” and ”귀밝이술” in Korean; this helps the model to identify faithful knowledge, consequently leading to the correct answer. However, in the case of English, not only are these cultural nuances lost in the English question, but the model misunderstood the question to asking about “Dan-o”, another Korean traditional holiday. Considering that the model is able to identify faithful knowledge from CS setting, we can see that even though the model has correct knowledge, activation is not successfully done on English setting. Solar lacks in knowledge about “정월대보름” in English, or fails to match its encoded knowledge with its English translation. We also provide a failure case of CS text in knowledge activation in Appendix C.3.
5.4 Error Analysis
@@
Although we informed the models that a correct answer is in one of the choices, we notice that the majority of incorrect responses were “None of the above”. The errors may derive from either hallucinated knowledge or failing to follow instructions faithfully. Therefore, we provide additional analysis on erroneous outputs in Table 9. We report the results in the format of number of None errors / the number of errors that derived from knowledge hallucination. Errors that are not from hallucination are caused by poor instruction-following.
LLMs tend to hallucinate much frequently when given English questions.
Overall, we observe that performance on English questions results in more errors compared to CS across all LLMs. This indicates that models tend to hallucinate much frequently when given English questions, again highlighting the effectiveness of CS over English setting. In particular, models that show excellent performance, such as GPT-4o, Claude 3.5, and Solar, generally yield less errors.222The results for GPT-4o in Society are an exception, where instruction-following errors consist most of the errors. It is also worth noting that Gemma2 families tend to hallucinate largely on the History and General domains, supporting our finding in Figure 4 and Figure 6 which respectively illustrates poor performance on human evaluation and QA accuracy.
@@ 6 Conclusion
In this paper, we explore the efficacy of code-switching in activating language-specific knowledge embedded in LLMs. Utilizing two Korean-centric QA datasets, we synthesize EnKoQA, a qualified English-Korean code-switching QA dataset. We formulate two tasks and conduct experiments across various multilingual LLMs. Our analyses demonstrate that code-switching can effectively activate knowledge within LLMs compared to English text, particularly in language-specific domains. Furthermore, we find that the performance gap between code-switching and English is more significant in models exhibiting strong Korean proficiency. Our work suggests the potential of code-switching as an effective strategy for eliciting language-specific knowledge from LLMs.
Limitation
@@
Another limitation of our work is that we conduct human evaluations on only a subset of LLMs, domains, and questions. Evaluating the quality (i.e., faithfulness and helpfulness) of knowledge in code-switched text presents inherent and practical challenges, as it necessitates evaluators to be fluent bilinguals. Consequently, we present only partial results for the knowledge identification task. To enhance the reliability of our evaluations for this task, we plan to employ the LLM-as-a-judge approach (Zheng et al., 2023), guided by the gold knowledge required for assessment.
Lastly, as we rely on a LLM, specifically gpt-3.5-turbo, to synthesize our code-switching dataset, the performance of the LLM can affect the quality of the dataset. To mitigate the risk of erroneous samples and to fully leverage the LLM’s capabilities, we engage reliable human annotators to review the samples and verify their quality.
@@ A.1 Details of Source Data
We used two datasets in our experiments:
- HAE RAE BENCH 1.1: This dataset is available at https://huggingface.co/datasets/HAERAE-HUB/HAE_RAE_BENCH_1.1.
- CLIcK: This dataset is available at https://huggingface.co/datasets/EunsuKim/CLIcK.
A.2 Statistics of EnKoQA
We provide statistics of EnKoQA per domain in Table 2.
@@ A.2 Statistics of EnKoQA
A.3 Quality Control Guideline
@@
A.4 Data Sample
@@ Appendix B Experimental Details
We used claude-3-5-sonnet API from Anthropic AI333https://www.anthropic.com/. The cost for claude-3-5-sonnet was $21 for experiment inference.
@@
We used solar-mini API from Upstage444https://www.upstage.ai/.

### v2

@@ 4.3 Experimental Setup
We conduct experiments with the same models as in Section 3.2. For knowledge identification, we instruct the model to write a list of factual knowledge that are required for solving the given question in one or two sentences. For knowledge leveraging, we pass on previously identified knowledge and ask the model to select an answer and explain why. The full-length prompts are provided in Table 14 and 15.
Evaluating Knowledge Identification.
In order to effectively evaluate knowledge identification results, we refer to Section 3.3 and choose two domains where CS performance is higher (i.e., History, Tradition), and two domains that have minimum difference (i.e., General, Law). Moreover, we select four models with different performances and sizes (i.e., GPT-4o, Solar, Gemma2 27B, Gemma2 9B). Specifically, we sample 10 questions from each domain and model, resulting in 160 samples. Then, we conduct human and LLM-based evaluation on identified knowledge.
Human Evaluation
We employ four human evaluators who are fluent in both Korean and English and completed Korean public education, thus qualified to evaluate questions sourced from Korean proficiency tests for foreigners and the Korean College Scholastic Ability Test. For faithfulness and helpfulness, the evaluator is asked to rate a knowledge list on a Likert scale from 1 to 3. In pairwise evaluation, we provide two knowledge lists in a random order and ask the evaluator to select a list that is overall more effective for answering the question. Details on evaluation criteria and evaluator information are provided in Appendix C.1 and C.2.
LLM-based Evaluation
As we conduct human evaluation on quite small amount of samples, we additionally conduct LLM-as-a-judge evaluation (Zheng et al., 2023) to amplify our analysis. Specifically, we use GPT-4o as the evaluator, using identical instructions with human evaluators on 40 questions for 9 domains and 8 models, 360 samples in total. Full prompts are provided in Appendix C.1.
5 Analysis on Knowledge Identification
5.1 Human Evaluation
@@
In the upper row of Figure 2, we observe a significant gap in faithfulness scores between CS and English in both History and Tradition. The discrepancy is more salient in Tradition where cultural nuances is much important, implying that asking questions in CS is much successful in capturing cultural nuances and meanings. In General domain, the scores for CS and English are almost identical (or even better in English for Gemma2 9B), indicating that the difference in knowledge activated by CS questions compared to English questions is minimal when addressing general and common facts. In Law, although knowledge from CS is slightly more faithful than that from English, their absolute scores are lower than those in other domains, suggesting that models fail to identify faithful knowledge that requires domain expertise.
@@
The lower row of Figure 2 presents evaluation results for helpfulness. It is intuitive that faithful knowledge serves as a valuable source for answering questions, and as a result, the evaluation of helpfulness shows a similar trend to that of faithfulness. In History and Tradition, the gap between CS and English becomes larger in helpfulness, emphasizing the effectiveness of the CS setting in identifying both faithful and helpful knowledge. It is also notable that the scores for helpfulness are particularly high for GPT-4o and Solar, models in which performance in CS surpasses that in English to a large extent (§ 3.3). In contrast, the helpfulness scores in the Law domain are considerably lower for both CS and English compared to other domains. Given that the Law domain requires expert-level legal knowledge, the models struggle to grasp the legal context, leading to difficulties in accurately identifying helpful knowledge sources from both CS and English questions.
@@
In Figure 3, the win ratio for CS is higher in History and Tradition, demonstrating that CS questions can activate more essential knowledge sources for question answering. On the contrary, in domains where CS does not show its effectiveness, the win ratio of CS is comparatively lower (i.e., General) or the ratio of Tie is high (i.e., Law). Especially in the case of Law, the quality of knowledge lists generated from CS questions is evaluated as equivalent to, or even worse than, that generated from English questions.
5.2 LLM-based Evaluation
We observe in Figure 6 and Figure 7 that the score gap between CS and English in both faithfulness and helpfulness are minimal. In fact, CS scores are even or lower for some cases, which are inconsistent with human evaluation results. However, it is still worth noting that LLM-as-a-judge also assigns higher scores for advanced models, and overall scores were lower in History and Tradition.
On the other hand, LLM judgement scores in pairwise evaluation generally agree with the human evaluations. We compute Cohen’s Kappa () score in Table 6, and follow interpretations from Landis and Koch (1977).222Landis and Koch (1977) interprets 0–0.20 as slight, 0.21–0.40 as fair, 0.41–0.60 as moderate, 0.61–0.80 as substantial, and 0.81–1 as almost perfect agreement. Consistent with human evaluation, the LLM judge votes CS for most cases, and the agreement is stronger with advanced models (i.e., GPT-4o), on culture-intensive domains (i.e., History, Tradition).
While other domains fairly agree with human judgment, Law shows exceptional results. Specifically, the LLM-as-judge evaluation reports a significantly higher win ratio for CS in the Law domain compared to human evaluation. However, considering that tie ratio is substantial in human evaluation as well, we speculate that LLM-as-a-judge gives a win to CS on knowledge that human evaluators regarded comparable quality with English setting.
6 Analysis on Knowledge Leveraging
We present the visualized results of accuracy in both CS and English settings in Figure 5, with detailed scores reported in Table 7.
Main Observations.
@@ 6 Analysis on Knowledge Leveraging
Consistent with the results in Section 3.3, all models demonstrate generally higher performances for CS questions compared to English questions. The results indicate that CS effectively activates knowledge across various domains while activating in dominant English language is suboptimal. GPT-4o, Claude, and Solar exhibit higher CS performance than English across all domains. These models not only identify faithful and helpful knowledge (§ 5.1), but also answer questions by accurately grounding on that knowledge; this shows that CS questions robustly activate essential knowledge in these models. On the contrary, Llama3 and Gemma2 families show poor performance in both CS and English settings in several domains, such as Geography and Law. Taking into account that these domains require domain-specific expertise, it is likely that their lack of understanding contributes to low accuracy, let alone CS failing to activate Korea-focused knowledge.
Knowledge Identification and Leveraging both matters.
We demonstrate that qualified knowledge identification is prerequisite for knowledge activation of CS. The win ratio of History knowledge by GPT-3.5 was relatively poor compared to others (Figure 4), leading GPT-3.5 to be the only model that did not benefit from CS. Similarly in Law, Figure 2 and 3 show that helpfulness and pairwise scores for knowledge by Gemma2 9B and 27B are lower than others, which are responsible for their suboptimal performance in CS. We provide further qualitative case study in Appendix D.4.
English questions hallucinate more than CS.
@@
Although we informed the models that the answer is in one of the choices, we notice that the majority of incorrect responses were “None of the above”. The errors may derive from either hallucinated knowledge or failing to follow instructions faithfully. Therefore, we provide additional analysis on erroneous outputs in Table 8. We report the results in the format of number of errors that derived from knowledge hallucination / total number of None errors. Errors that are not from hallucination are caused by poor instruction-following. Overall, we observe that performance on English questions results in more errors compared to CS across all LLMs, and most of them were hallucination errors. This indicates that models hallucinate much frequently when English questions are given, again highlighting the effectiveness of CS over English setting. It is also worth noting that Gemma2 families hallucinate largely on History and General, supporting our finding in Figure 2 and Figure 5 which respectively illustrates poor performance on human evaluation and QA accuracy.
@@ 7 Conclusion
We explore the efficacy of code-switching in activating language-specific knowledge embedded in LLMs. Utilizing two Korean-centric QA datasets, we synthesize EnKoQA, a qualified English-Korean code-switching QA dataset and conduct experiments on various multilingual LLMs. Our analyses demonstrate that LLMs can simulate a similar code-switching effect with human communications of facilitating low-resource knowledge within LLMs, particularly in language-specific domains. Regarding this finding, we suggest that code-switching can be an effective strategy for solving low-resource language tasks. Also, augmenting low-resource datasets into code-switching text can amplify resource and mitigate data scarcity challenge. We hope our work can motivate NLP community to explore more potential of code-switching for developing robust multilingual LLMs.
Limitations and Future Work
@@
Another limitation of our work is that we conduct human evaluations on only a subset of LLMs, domains, and questions. Evaluating the quality (i.e., faithfulness and helpfulness) of knowledge in code-switched text presents inherent and practical challenges, as it necessitates evaluators to be fluent bilinguals. Consequently, we present only partial results for the knowledge identification task.
Lastly, as we rely on a LLM, specifically gpt-3.5-turbo, to synthesize our code-switching dataset, the performance of the LLM can affect the quality of the dataset. To mitigate the risk of erroneous samples and to fully leverage the LLM’s capabilities, we engage reliable human annotators to review the samples and verify their quality. Also, as we formulate our code-switching dataset with gold English and Korean, in a more realistic scenario where a monolingual English speaker creates code-switching text, sentences would be created automatically without any additional supervision.
In the future, we aim to investigate more potential of code-switching in diverse aspects, including instruction-tuning of LLMs to users effectively using code-switching for multilingual tasks. As we have demonstrated synthesizing monolingual datasets into code-switching text, we hope our work can inspire NLP community to explore the capability of code-switching in enhancing and utilizing multilingual LLMs.
@@ A.1 Details of Source Data
CLIcK (Kim et al., 2024b)333https://huggingface.co/datasets/EunsuKim/CLIcK consists of 1,995 multiple-choice QA pairs, classified in two main categories (Culture, Language) and 11 sub-categories. CLIcK is sourced from various official Korean exams and textbooks, e.g., College Scholastic Ability Test of Korea (CSAT). In this work, we only utilize data of eight sub-categories from Korean Culture category as our work aims to evaluate the effect of CS on activating Korean-specific knowledge.
HAE-RAE (Son et al., 2024)444https://huggingface.co/datasets/HAERAE-HUB/HAE_RAE_BENCH_1.1 is a Korean benchmark dataset originally crafted to capture cultural and contextual nuances inherent to the Korean language. We use 1,027 multiple-choice QA pairs regarding Korean culture. Both datasets are sourced from official Korean exams, textbooks, and text on the internet.
We combine two datasets and merge common categories (i.e., Society, Geography, and Law), resulting in 2,372 QA pairs in nine categories: Popular, Economy, Politics, Tradition, General Knowledge, Society, Geography, History, and Law.
A.2 Dataset Statistics and License
We provide statistics of EnKoQA per domain in Table 2. We plan to release the dataset in public, under CC BY-NC license. We clarify that the source datasets are either open-source or used under authors’ permission, ensuring that there are no issues regarding their use.
@@ A.2 Dataset Statistics and License
A.3 Code-Mixing Index
| CMI | Tradition | History | ||||||
| Solar | Gemma 2 9B | Gemma 2 27B | GPT-4o | Solar | Gemma2 9B | Gemma2 27B | GPT-4o | |
| 0–10 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 |
| 10–20 | 100.0 | 100.0 | 100.0 | 100.0 | 50.00 | 50.00 | 100.0 | 50.00 |
| 20–30 | 56.25 | 65.62 | 53.12 | 71.88 | 50.00 | 31.58 | 36.84 | 60.53 |
| 30–40 | 72.34 | 59.57 | 55.32 | 76.6 | 66.67 | 48.72 | 45.3 | 77.78 |
| 40–50 | 80.39 | 62.75 | 70.59 | 84.31 | 69.54 | 42.38 | 42.38 | 78.15 |
| 50–60 | 83.72 | 74.42 | 74.42 | 93.02 | 62.35 | 40.00 | 35.29 | 74.12 |
| 60–70 | 85.19 | 66.67 | 66.67 | 92.59 | 50.00 | 28.85 | 46.15 | 63.46 |
| 70–80 | 66.67 | 58.33 | 66.67 | 91.67 | 57.89 | 15.79 | 21.05 | 57.89 |
| 80–90 | 66.67 | 83.33 | 83.33 | 100.0 | 50.00 | 50.00 | 25.00 | 100.0 |
| 90–100 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 | 00.00 |
We report CMI scores for our dataset in Table 3. In specific, we tokenized the sentence using bert-base-multilingual-cased, then removed all noisy tokens such as numbers or tags and counted the ratio of . We report the distribution of QA accuracy on different CMI scores in Tradition and History, two domains where CS proved its effectiveness. If CMI is close to 0, sentence is mostly written in English, and close to 100 means vice versa. The number of samples at each end (0-10, 90-100) was very small, causing outliers. We can see that accuracy scores are quite evenly distributed across all ratios, suggesting that there is no distinct tendency between CMI and accuracy performance.
It is important to note, however, that code switching metrics such as CMI, while offering a quantitative measure of token-level composition, are inherently limited in capturing the nuanced semantic and syntactic characteristics of code-switched texts. These metrics primarily rely on surface-level token ratios, which can inadvertently assign high scores to linguistically or contextually meaningless sequences. Consequently, they may over-represent the presence of meaningful code-switching patterns while failing to account for the deeper linguistic interplay that defines effective code-switching. For a more comprehensive discussion of these limitations, please refer to Srivastava and Singh, 2021.
A.4 Quality Control Guideline
@@
A.5 Annotation Details
For dataset construction, two Korean native annotators with expert knowledge in Korean culture and equivalently fluent in English manually examine the candidates and select the most naturally code-switched question, then cross-checked each other’s assigned share of dataset. If a selected candidate appeared to be incorrect or suboptimal, the annotators engaged in thorough discussions until they reached an agreement on the most appropriate candidate.
Regarding inter-annotator agreement (IAA), although we did not compute a formal IAA score, significant effort was devoted to ensuring high annotation quality through extensive discussion and collaboration among annotators. In specific, the annotation process involved annotators who are fluent in both English and Korean are assigned each portion of the dataset to select a candidate for code-switched question. Following this initial annotation, the annotators cross-checked each other’s work to identify any discrepancies. If a selected candidate appeared to be incorrect or suboptimal, the annotators engaged in thorough discussions until they reached an agreement on the most appropriate candidate. This iterative and collaborative process was integral to constructing a high-quality dataset.
A.6 Dataset Size and Quality
Discussion on Dataset Size
While we acknowledge the relatively limited size of EnKoQA dataset, we emphasize that quality often matters more than quantity as many studies (Pacchiardi et al., 2024; Maia Polo et al., 2024; Vivek et al., 2024) have demonstrated. Please note that we prioritized creating a high-quality dataset with rigorous manual validation and linguistic alignment, ensuring that the dataset serves as a reliable resource for code-switching research. Additionally, while the size of Korean datasets is often limited given that Korean is a low-resource language, EnKoQA dataset is comparatively larger than the sizes of other Korean datasets. For instance, datasets in the Open Ko-LLM leaderboard (Park et al., 2024), such as Ko-ARC (1.1k), Ko-TruthfulQA (0.8k), and Ko-CommonGen (0.8k), are all smaller in scale than EnKoQA’s 2,372 question-answer pairs. This highlights our effort to provide a relatively extensive resource within the constraints of dataset availability for minor languages.
Specifically, our quality control process includes human annotators thoroughly reviewing all LLM-generated samples to assess the quality and naturalness. When any errors or unnatural code-switching patterns were identified, annotators corrected them to ensure that the final dataset adheres to high standards of our quality control. In that sense, GPT-3.5-turbo served as an assistive tool for providing initial candidates, rather than generating final outputs. Therefore, we assert that any potential shortcomings of the translation tool were effectively mitigated through this meticulous human review and correction process.
@@ A.6 Dataset Size and Quality
Translating with GPT-3.5
We have conducted experiments on both GPT-3.5 and GPT-4o for translation and code-switching generation tasks. Interestingly, we observed that after manual examination and correction process, the results from both models were comparable in terms of quality and naturalness. This is due to our rigorous human-in-the-loop workflow that ensures any errors or unnatural expressions are taken care of, regardless of the initial model used. Given this finding, we used GPT-3.5 for its cost efficiency while maintaining high-quality standards through meticulous human examination and refinement. By prioritizing manual validation, we ensured that the final dataset reflects linguistic accuracy and naturalness, independent of the model used for preliminary generation.
A.7 Data Sample
@@ Appendix B Experimental Details
We used claude-3-5-sonnet API from Anthropic AI555https://www.anthropic.com/. The cost for claude-3-5-sonnet was $21 for experiment inference.
@@
We used solar-mini API from Upstage666https://www.upstage.ai/.